### v2

# We prove that if $A=\{a_{1},\dots,a_{|A|}\}\subset\{1,2,\dots,n\}$ is a Sidon set so that $|A|=n^{1/2}-L^{\prime}$, then

Abstract
We prove that if $A=\{a_{1},\dots,a_{|A|}\}\subset\{1,2,\dots,n\}$ is a Sidon set so that $|A|=n^{1/2}-L^{\prime}$, then
$a_{m}=m\cdot n^{1/2}+\mathcal{O}\left(n^{7/8}\right)+\mathcal{O}\left(L^{1/2}\cdot n^{3/4}\right)$
where $L=\max\{0,L^{\prime}\}$. As an application of this, we give easy proofs of some previously derived results. We proceed on to proving that for any $\varepsilon>0$, we have
$\sum_{a\in S}a=\frac{1}{2}n^{3/2}+\mathcal{O}\left(n^{11/8}\right)$
for all $n\leq N$ but at most $\mathcal{O}_{\varepsilon}\left(N^{\frac{3}{5}+\varepsilon}\right)$ exceptions.
The $m$-th Element of a Sidon Set
R. Balasubramanian
The Institute of Mathematical Sciences, Chennai
balu@imsc.res.in
Sayan Dutta
Department of Mathematics and Statistics, IISER Kolkata, Mohanpur, West Bengal 741246, India
sd19ms148@iiserkol.ac.in
key-word: Sidon set
1 Introduction
A set of positive integers $A\subset\mathbb{N}$ is called a Sidon Set or a Sidon Sequence if the equation $a+b=c+d$ does not have any non-trivial solutions in $A$. They were named after Hungarian mathematician Simon Sidon who was inspired by certain problems in Fourier series to ask Erdős about the possible growth of such sequences.
Since then, there has been an extensive amount of literature on this topic exploring a plethora of different questions about finite and infinite Sidon sets [26]. Addressing the original question of Sidon, Erdős conjectured [13] (and offered \$ 500 for a proof or disproof) that if $S(n)$ denotes the maximum possible cardinality of a Sidon subset of $[n]:=\{1,2,\dots,n\}$, then
$S(n)<n^{1/2}+o(n^{\varepsilon})$
for all $\varepsilon>0$.
Several attempts on this problem by several different authors have only yielded
$S(n)\leq n^{1/2}+\mathcal{O}\left(n^{1/4}\right)$
although a variety of different techniques have been tried [24, 31, 6, 16]. A classic result by Bose and Chowla [3] shows that the conjectured bound cannot be improved.
The most recent result on this was provided by Balogh et. al. [1] proving
$S(n)<n^{1/2}+(1-\gamma)n^{1/4}$ (1)
for some $\gamma\geq 0.002$. In the same spirit, Kevin O’Bryant [27] improved the constant from $0.998$ to $0.99703$, and then again to $0.98183$ with Daniel Carter and Zach Hunter [4].
Another recent improvement in the theory of Sidon sets is a solution to the Erdős Sárközy Sós problem on asymptotic Sidon bases of order $3$. After a series of developments made by Deshoulliers and Plagne [9], Kiss [22], Kiss, Rozgonyi and Sándor [23] and Cilleruelo [8], it was finally solved by Cédric Pilatte [29] using some recent results of Will Sawin [33].
However our focus is on finding a formula for the $m$-th element of a finite Sidon set. As an application of our formula, we will indicate easier proofs of the theorems recently obtained by Yuchen Ding [10, 11].
Our main tool is a theorem due to Cilleruelo.
Theorem A (Cilleruelo [5]).
Let $A\subset[n]$ be a Sidon set with $\left\lvert A\right\rvert=n^{1/2}-L^{\prime}$. Then, every subinterval $I\subset[1,n]$ with length $cn$ contains $c\left\lvert A\right\rvert+E_{I}$ elements of $A$ where
$E_{I}\leq 52\;n^{1/4}\left(1+c^{1/2}\cdot n^{1/8}\right)\left(1+L^{1/2}\cdot n^{-1/8}\right)$
where $L=\max\{0,L^{\prime}\}$.
2 Dense Sidon Sets
Definition 1.
A Sidon subset $A\subset[n]$ is called dense if
$\left\lvert A\right\rvert=\max\left\lvert S\right\rvert$
where the maximum is taken over all Sidon subsets of $[n]$.
The structure of dense Sidon sets has a rich literature [12, 30] and classic constructions by Erdős–Turán [14], Singer [34], Bose [2], Spence [15, 31], Hughes [21] and Cilleruelo [7] have established that a dense Sidon set $A$ satisfies $\left\lvert A\right\rvert\geq\left(1-o(1)\right)\sqrt{n}$. As remarked by Ruzsa, “somehow all known constructions of dense Sidon sets involve the primes” [32].
We will begin by proving a lemma that will justify an assumption we will make in the results to follow. This requires a classic result of Bose and a recent result of Baker, Harman and Pintz.
Theorem B (R. C. Bose [2]).
For a prime $p$, there are at least $p$ elements in $[p^{2}-1]$ such that the sums of two of these elements are different modulo $p^{2}-1$.
Theorem C (Glyn Harman, R.C. Baker, J. Pintz. [17]).
We have
$p_{k+1}-p_{k}\ll p_{k}^{0.525}$
where $p_{k}$ is the $k$-th prime.
Lemma 2.
Let $A\subset[n]$ be a dense Sidon set so that $|A|=n^{1/2}-L$. Then $-n^{1/4}\leq L\leq n^{21/80}$.
Proof.
As discussed earlier, from Equation (1), we immediately have $L\geq-n^{1/4}$.
For the next part, let us denote
$S_{n}=\max\left\lvert S\right\rvert$
over all Sidon subsets $S\subset[n]$.
Also assume
$p_{k}^{2}-1<n\leq p_{k+1}^{2}-1$
where $p_{i}$ is the $i$-th prime. This gives
$p_{k}\leq S_{n}$
using Theorem B.
But from Theorem C,
$p_{k+1}-p_{k}\ll p_{k}^{21/40}\leq n^{21/80}$
hence giving
$p_{k}\geq\sqrt{n}-n^{21/80}$
thus completing the proof.
Remark: Notice that if we assume $p_{k+1}-p_{k}\ll\sqrt{p_{k}}$, then an exact same line of argument will give $L\leq n^{1/4}$.
3 The $m$-th element
As advertised, we will now give the main theorem of the paper.
Theorem 3.
Let $A=\{a_{1},\dots,a_{\left\lvert A\right\rvert}\}\subset[n]$ be a dense Sidon set with $|A|=n^{1/2}-L^{\prime}$. Then
$a_{m}=m\cdot n^{1/2}+\mathcal{O}\left(n^{7/8}\right)+\mathcal{O}\left(L^{1/2}\cdot n^{3/4}\right)$
where $L=\max\{0,L^{\prime}\}$.
Proof.
Consider a Sidon set $A=\{a_{1},\dots,a_{|A|}\}\subset[n]$ so that $|A|=n^{1/2}-L$ with $L\leq n^{\frac{21}{80}}$. Let $I=[a_{m}]\subset[n]$ Let $a_{m}=cn+1$ so that $c=\dfrac{a_{m}-1}{n}<1$. So, $|I|=cn$ and $\left\lvert A\cap[a_{m}]\right\rvert=m$.
By Theorem A, the number of elements in an interval $I$ of length $cn$ is
$c\left\lvert A\right\rvert+\mathcal{O}\left(n^{1/4}\right)+\mathcal{O}\left(n^{3/8}\right)+\mathcal{O}\left(L^{1/2}\cdot n^{1/8}\right)+\mathcal{O}\left(L^{1/2}\cdot n^{1/4}\right)$
for $c\leq 1$.
This gives
$m$ $=c\left(n^{1/2}+\mathcal{O}\left(L\right)\right)+\mathcal{O}\left(n^{3/8}\right)+\mathcal{O}\left(L^{1/2}\cdot n^{1/4}\right)$
$=cn^{1/2}+\mathcal{O}\left(L\right)+\mathcal{O}\left(n^{3/8}\right)+\mathcal{O}\left(L^{1/2}\cdot n^{1/4}\right)$
and hence putting $c=\dfrac{a_{m}-1}{n}$, multiplying by $n^{1/2}$, and rearranging, we get
$a_{m}=m\cdot n^{1/2}+\mathcal{O}\left(n^{7/8}\right)+\mathcal{O}\left(L^{1/2}\cdot n^{3/4}\right)$
thus completing the proof.
This gives us the immediate corollary.
Corollary 4.
Let $A=\{a_{1},\dots,a_{|A|}\}\subset[n]$ be a Sidon set so that $|A|=n^{1/2}-L^{\prime}$ with $L\leq n^{\frac{21}{80}}$. Then,
$a_{m}^{\ell}=m^{\ell}\cdot n^{\frac{\ell}{2}}+\mathcal{O}\left(m^{\ell-1}\cdot n^{\frac{4\ell+3}{8}}\right)+\mathcal{O}\left(m^{\ell-1}\cdot L^{1/2}\cdot n^{\frac{2\ell+1}{4}}\right)+\mathcal{O}\left(n^{\frac{7\ell}{8}}\right)+\mathcal{O}\left(L^{\frac{\ell}{2}}\cdot n^{\frac{3\ell}{4}}\right)$
for any positive integer $\ell$.
Proof.
This follows from the more general statement that if
$f(x)=m(x)+\mathcal{O}\left(e(x)\right)$
then
$\left(f(x)\right)^{\ell}=\left(m(x)\right)^{\ell}+\mathcal{O}\left(e(x)\left(m(x)\right)^{\ell-1}\right)+\mathcal{O}\left(e(x)^{\ell}\right)$
using the Binomial Theorem.
We now also have the following corollaries that were also derived in [10] and [11].
Corollary 5.
Let $A=\{a_{1},\dots,a_{|A|}\}\subset[n]$ be a Sidon set so that $|A|=n^{1/2}-L^{\prime}$. Then,
$\sum_{a\in A}a=\frac{1}{2}\cdot n^{3/2}+\mathcal{O}\left(n^{\frac{11}{8}}\right)+\mathcal{O}\left(L^{1/2}\cdot n^{5/4}\right)$
for $L\leq n^{\frac{21}{80}}$.
Proof.
Using Theorem 3 and using the fact that $\left\lvert A\right\rvert<2\sqrt{n}$, we have
$\sum_{m=1}^{\left\lvert A\right\rvert}a_{m}$ $=\frac{\left\lvert A\right\rvert\left(\left\lvert A\right\rvert+1\right)}{2}\sqrt{n}+\mathcal{O}\left(n^{7/8}\left\lvert A\right\rvert\right)+\mathcal{O}\left(L^{1/2}n^{3/4}\left\lvert A\right\rvert\right)$
$=\frac{\sqrt{n}}{2}\left(\sqrt{n}-L\right)^{2}+\frac{\sqrt{n}}{2}\left(\sqrt{n}-L\right)+\mathcal{O}\left(n^{7/8}\sqrt{n}\right)+\mathcal{O}\left(L^{1/2}\cdot n^{3/4}\sqrt{n}\right)$
$=\frac{\sqrt{n}}{2}\left(n+L^{2}-2L\sqrt{n}\right)+\frac{\sqrt{n}}{2}\left(\sqrt{n}-L\right)+\mathcal{O}\left(n^{\frac{11}{8}}\right)+\mathcal{O}\left(Ln^{7/8}\right)+\mathcal{O}\left(L^{1/2}n^{5/4}\right)+\mathcal{O}\left(L^{3/2}n^{3/4}\right)$
$=\frac{1}{2}\cdot n^{3/2}+\mathcal{O}\left(n^{\frac{11}{8}}\right)+\mathcal{O}\left(L^{1/2}\cdot n^{5/4}\right)$
hence completing the proof.
Corollary 6.
Let $A=\{a_{1},\dots,a_{|A|}\}\subset[n]$ be a Sidon set so that $|A|=n^{1/2}-L^{\prime}$. Then,
$\sum_{a\in A}a^{\ell}=\frac{1}{\ell+1}\cdot n^{\frac{2\ell+1}{2}}+\mathcal{O}\left(n^{\frac{8\ell+3}{8}}\right)+\mathcal{O}\left(L^{1/2}\cdot n^{\frac{4\ell+1}{4}}\right)$
for $L\leq n^{\frac{21}{80}}$.
Proof.
Using Corollary 4 in exactly the same way as in the previous proof, we get
$\sum_{m=1}^{\left\lvert A\right\rvert}a_{m}^{\ell}$ $=n^{\frac{\ell}{2}}\cdot\sum_{m=1}^{\left\lvert A\right\rvert}m^{\ell}+\mathcal{O}\left(n^{\frac{4\ell+3}{8}}\cdot\sum_{m=1}^{\left\lvert A\right\rvert}m^{\ell-1}\right)+\mathcal{O}\left(L^{1/2}\cdot n^{\frac{2\ell+1}{4}}\cdot\sum_{m=1}^{\left\lvert A\right\rvert}m^{\ell-1}\right)$
$\quad\quad\quad\quad\quad\quad\,+\mathcal{O}\left(n^{\frac{7\ell}{8}}\cdot\sqrt{n}\right)+\mathcal{O}\left(L^{\frac{\ell}{2}}\cdot n^{\frac{3\ell}{4}}\cdot\sqrt{n}\right)$
$=\frac{1}{\ell+1}\left(\sqrt{n}\right)^{\ell+1}\cdot n^{\frac{\ell}{2}}+\mathcal{O}\left(\left(\sqrt{n}\right)^{\ell}\cdot n^{\frac{4\ell+1}{8}}\right)+\mathcal{O}\left(\left(\sqrt{n}\right)^{\ell}\cdot L^{1/2}\cdot n^{\frac{2\ell+1}{4}}\right)$
$=\frac{1}{\ell+1}\cdot n^{\frac{2\ell+1}{2}}+\mathcal{O}\left(n^{\frac{8\ell+3}{8}}\right)+\mathcal{O}\left(L^{1/2}\cdot n^{\frac{4\ell+1}{4}}\right)$
hence completing the proof.
Remark: It should be noted that Theorem 3 and Corollary 4 are valid formulas for all $m$. However, they are only useful as an asymptotic formula when $m$ is close to $\sqrt{n}$.
4 Improvements
We will now provide an improvement to Theorem 1.4 and Corollary 1.5 of Ding [11]. We will require the following result due to Heath-Brown which is an improvement on a series of similar results by Wolke [35], Heath-Brown [18, 19], Peck [28] and Matomäki [25].
Theorem D (Roger Heath-Brown [20]).
For any $\varepsilon>0$ we have
$\sum_{\begin{subarray}{c}p_{n}\leq x\\ p_{n+1}-p_{n}\geq\sqrt{p_{n}}\end{subarray}}\left(p_{n+1}-p_{n}\right)\ll_{\varepsilon}x^{\frac{3}{5}+\varepsilon}$
where $p_{n}$ is the $n$-th prime.
We finally express
Theorem 7.
Let $S$ be a dense Sidon set in $[n]$. Then, for any $\varepsilon>0$, we have
$\sum_{a\in S}a=\frac{1}{2}n^{3/2}+\mathcal{O}\left(n^{\frac{11}{8}}\right)$
for all $n\leq N$ but at most $\mathcal{O}_{\varepsilon}\left(N^{\frac{3}{5}+\varepsilon}\right)$ exceptions.
Proof.
We have already proven that
$\sum_{a\in S}a=\frac{1}{2}\cdot n^{3/2}+\mathcal{O}\left(n^{\frac{11}{8}}\right)+\mathcal{O}\left(L^{1/2}\cdot n^{5/4}\right)$
for any choice of $L$.
So, it is enough to show that $\left\lvert L\right\rvert<n^{1/4}$ happens almost always with a small number of exceptions. As noticed in the remark under Lemma 2, denoting $p_{m}$ to be the $m$-th prime, if we have
$p_{m+1}-p_{m}<\sqrt{p_{m}}$
then for all $n\in\left[p_{m},p_{m+1}\right]$, we have $\left\lvert L\right\rvert<n^{1/4}$.
So, the set of exceptions are precisely those $n\in\left[p_{m},p_{m+1}\right]$ with $p_{m+1}-p_{m}\geq\sqrt{p_{m}}$. The number of such $n$ is given by
$\sum_{\begin{subarray}{c}p_{n}\leq N\\ p_{n+1}-p_{n}\geq\sqrt{p_{n}}\end{subarray}}\left(p_{n+1}-p_{n}\right)\ll_{\varepsilon}N^{\frac{3}{5}+\varepsilon}$
which follows from Theorem D.
This completes our proof.
Remark: An exact same argument gives
$\sum_{a\in A}a^{\ell}=\frac{1}{\ell+1}\cdot n^{\frac{2\ell+1}{2}}+\mathcal{O}\left(n^{\frac{8\ell+3}{8}}\right)$
for all $n\leq N$ but at most $\mathcal{O}_{\varepsilon}\left(N^{\frac{3}{5}+\varepsilon}\right)$ exceptions.
We conclude by making an observation on Theorem 1.2 of Ding [10]. It essentially says that the sum of elements of a Sidon set $A$ in the residue class $i\;(\mathrm{mod}\;m)$ is asymptotically $\frac{1}{2m}n^{3/2}$ provided that there is a function $f(n)\to\infty$ as $n\to\infty$ such that for all $t\in\left(\frac{n}{f(n)},n\right)$, we have $A(t):=A\cap(0,t)>\sqrt{t}$. However, it seems that this condition will rarely be satisfied for a dense Sidon set. For example, let us take $t=0.01n$. By Theorem A, $A(t)\sim 0.01\sqrt{n}<0.1\sqrt{n}=\sqrt{t}$. In other words, Theorem A prevents the said condition from being true under the assumption that $A(n)>\sqrt{n}$.
5 Acknowledgements
We would like to thank Prof. Kevin O’Bryant for bringing [27] and [4] to our attention.